A wire-wound resistor with a resistance of 250ohm will overheat if the power dissipated exceeds 10W. What is the maximum constant voltage V that can be applied across the terminals of this resistor?
1
Expert's answer
2020-04-23T12:45:27-0400

"P=IU, I=\frac {U}{R}
 \implies\\
P=\frac{U^2}{R}\implies
U=\sqrt{PR}\implies\\
U=\sqrt{10\cdot250}=\sqrt{2500}=50 \textnormal{v}\\
answer: U=50\textnormal{volt}"